The photography club has 28 members. There are 12 boys in the club.
What is the ratio of boys to girls in simplest form?

Answer by stanbon(75887)

The photography club has 28 members. There are 12 boys in the club.
What is the ratio of boys to girls in simplest form?
# of girls = 28-12 = 16
---
Ans:: 16/12 = 4 to 3
----------------
Cheers,
Stan H.
--------------